# SVP: Style-Enhanced Vivid Portrait Talking Head Diffusion Model

### v1

@@ Abstract
Introduction
Recent advancements in generative models have shed light on generating high-quality and realistic videos under various controlling conditions such as texts (Rombach et al. 2022), images (Blattmann et al. 2023), videos (Wu et al. 2023), etc. Among all the different subtasks of video generation, Talking Head Generation (THG), as a human-centric task which aims to generate videos of talking heads guided by conditions such as speech and images, has emerged as a significant problem due to its wide application in scenarios such as digital humans, film production, virtual reality. In spite of its importance, this is one of the most challenging tasks in video generation, resulted from its low tolerance to artifacts in general and its demand of high fidelity in lip shapes, facial expressions, and head motions.
Following the commonly used generative models, the GAN-based THG methods (Prajwal et al. 2020; Zhang et al. 2023) have achieved remarkable results in generating high-resolution videos through adversarial training between generators and discriminators, particularly in terms of visual quality and lip-sync accuracy. Diffusion model-based THG methods (Shen et al. 2023; Tian et al. 2024), on the other hand, excel in generating high-quality and high-resolution images and videos, and it outperforms GANs in terms of the stability and consistency of the generated content, thus becoming the mainstream methods for THG. These methods largely facilitate THG by strengthening the explicit controlling conditions such as facial keypoints and head motion sequences. However, they generally ignore the important fact of talking head videos. Essentially, when different people present speeches in real-life cases, they could have significant differences in habits and emotions under various circumstances. Such a fact in turn leads to different attributes in the corresponding talking head videos, including the visemes and expressions. Consequently, these habits and emotions are embedded as the intrinsic style in talking head videos. This intrinsic style, while being highly related to whether a video is realistic, can hardly be inferred from conditions such as facial keypoints which are widely adopted by the previous methods. As a result, when there is a large gap in the intrinsic style between the reference face and the speaker of the style reference video, the previous methods struggle to reproduce the real situation accurately.
@@ Introduction
Specifically, our $\mathtt{SVP}$ focuses on two main problems, i.e. extracting intrinsic style embeddings from style reference videos and controlling diffusion models with such embeddings. For intrinsic style extraction, a naive solution would be following StyleTalk (Ma et al. 2023), which maps 3D Morphable Model (3DMM) (Blanz and Vetter 1999) expression coefficients of the style reference video to style-related features. However, since attributes like visemes and expressions vary along the video frames, the deterministic embedding would suffer from insufficient capacity to model the latent manifold of intrinsic styles. Moreover, as one of the main parts of the video, the use of corresponding audio, which contains abundant information regarding the intrinsic styles, was not explored in StyleTalk, leading to unrepresentative style embeddings.
@@
To validate the effectiveness of our proposed method, we conduct extensive experiments and comparisons on the MEAD and HDTF datasets. Our method significantly outperforms other competitors across multiple metrics, including FVD, FID, PSNR, and SSIM. In addition to quantitative evaluation, we also perform comprehensive qualitative assessments. The results indicate that our method can generate highly natural and expressive talking videos, and can produce different emotions or even multiple changes in expressions within the same video according to user needs, achieving satisfactory visual effects.
@@
  • We propose an intrinsic style extractor that captures and expresses the intrinsic style of individuals through a self-supervised approach. We also incorporate audio information as an auxiliary into the intrinsic style extractor, leveraging subtle variations and emotional information to enhance the intrinsic style features. This allows the model to more accurately reflect the emotions and habits of speakers.
  • We design a resampling method suitable for diffusion models, enhancing the stability and generalization.
Related Work
@@
There has been significant research on GAN-based methods for person-generic audio-driven talking head generation. Early methods (Prajwal et al. 2020; Cheng et al. 2022; Wang et al. 2023) achieved lip synchronization by establishing a discriminator that correlates audio with lip movements. Other approaches (Zhou et al. 2020, 2021; Wang et al. 2021, 2022; Zhang et al. 2023) generated portrait videos by mapping audio to key facial information, such as landmarks, key points, or 3D Morphable Model (3DMM) (Blanz and Vetter 1999) coefficients, before rendering the final frame. However, due to the limitations of GANs in terms of generative capacity, the results produced by these methods often suffer from artifacts like pseudo-textures or restricted motion ranges.
@@
Recently, there has been a surge of research  (Shen et al. 2023; Xie et al. 2024; Tian et al. 2024; Wang et al. 2024; Wei, Yang, and Wang 2024; Xu et al. 2024; Yang et al. 2024) utilizing diffusion models to achieve high-quality portrait videos. Among these, X-Portrait (Xie et al. 2024) and MegActor (Yang et al. 2024) rely on the pose and expression from the source video to generate the target video, which limits their ability to produce videos based solely on audio. DiffTalk (Shen et al. 2023) was the first to modify lip movements using audio in conjunction with diffusion models, but it does not extend to driving other head parts. EMO (Tian et al. 2024) was the first to leverage LDM (Rombach et al. 2022) and audio features to achieve overall motion in portraits. V-Express (Wang et al. 2024) controls the overall motion amplitude by adjusting audio attention weights, while Hallo (Xu et al. 2024) designed a hierarchical module to regulate the motion amplitude of different regions. In summary, current audio-driven diffusion model approaches have not taken into account that each portrait should exhibit a corresponding style while speaking, which is essential for generating higher-quality portrait videos.
@@
Previous research has explored several GAN-based methods  (Wang et al. 2020; Wu et al. 2021; Ji et al. 2021, 2022; Liang et al. 2022; Sinha et al. 2022; Ma et al. 2023) for extracting style information to apply in talking head generation.  (Wang et al. 2020) and  (Sinha et al. 2022) directly inject labels into the network to drive the corresponding emotions.  (Liang et al. 2022) and  (Ji et al. 2022) map the facial expressions of each frame in the source video to each frame in the target video.  (Wu et al. 2021) and  (Ma et al. 2023) employ 3D Morphable Models (3DMM) to extract facial information and construct style codes that drive the desired styles. Building on these approaches, our framework is the first to propose the extraction of intrinsic style priors by integrating audio and 3D facial information. Additionally, we design a probabilistic learning to enhance style control within diffusion models.
Method
@@
The goal of THG is to generate a talking head video under the control of a reference image, audio, Head-Kps image sequence, and intrinsic style prior. Among these conditions, the reference image controls the background and facial identity. The audio guides the lip movements. Each Head-Kps image controls the head pose for each generated frame. The Head-Kps images are synthesized by mapping 8 facial keypoints (i.e., left and right edges, the pupils of both eyes, and the bridge of the nose) onto a black background. In addition, we propose the probabilistic style prior learning to extract the style prior from a style reference video, which is used to determine facial emotion and speaking habits.
@@
In $\mathtt{SVP}$, we employ a Latent Diffusion Model (LDM) (Rombach et al. 2022) to generate video frames. The LDM uses a diffusion and denoising process in the latent space via a Variational Autoencoder (VAE). It maps the input image $x$ to the latent space, encoding the image as $z=E(x)$, which helps maintain visual quality while reducing computational cost. During the diffusion process, Gaussian noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is gradually introduced into the latent $z$, degrading it into complete noise $z\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ after $T$ steps. In the reverse denoising process, the target latent $z$ is iteratively denoised from the sampled Gaussian noise using the diffusion model and then decoded by the VAE decoder $D$ into the output image $x=D(z)$. During training, given the latent $z_{0}=E(x_{0})$ and condition $c$, the denoising loss is:
@@
Among them, $z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$ represents the noisy latent variables at timestep $t\in[1,T]$, and $\epsilon_{t}$ is the added noise. $\epsilon_{\theta}$ is the noise predicted by the UNet model, modified using an attention mechanism with parameters $\theta$. This model employs a cross-attention mechanism to fuse the condition cc with the latent features $z_{t}$ , thereby guiding the image generation. $\mathtt{SVP}$ uses Stable Diffusion v1.5 (SDv1.5), a text-to-image Latent Diffusion Model (LDM), as the backbone. SDv1.5 is implemented based on U-Net (Ronneberger, Fischer, and Brox 2015), with each Transformer (Vaswani 2017) block containing both self-attention and cross-attention layers.
@@
As depicted in Figure 2, Our $\mathtt{SVP}$ consists of two important designs, namely Probabilistic Style Prior Learning and Style-Driven Diffusion Process. In Probabilistic Style Prior Learning, we propose the style extractor takes the 3DMM expression parameters and audio features as inputs to generate the style prior, which is represented as a Gaussian distribution. In the Style-Driven Diffusion Process, the reference image, Head-Kps sequences, audio features, and the style prior are input as control conditions into the Denoising UNet through Head-Kps Guider and attention layers. Finally, the vivid portrait frames are generated by the VAE decoder after the iteration of the denoising process.
Probabilistic Style Prior Learning
In order to learn representative intrinsic style indicators from style reference videos, we propose the novel probabilistic style prior learning. Built upon the transformer-based style encoder as in StyleTalk (Ma et al. 2023), we adopt a novel framework to make better usage of the style-related information contained in each video. Concretely, for a video clip, we first transform it into its corresponding frame-level audio parameters $\alpha\in\mathbb{R}^{N\times 1920}$ via Whisper-Tiny (Radford et al. 2023) and sequential expression parameters $\beta\in{\mathbb{R}}^{N\times 64}$ via the 3DMM encoder, where $N$ denotes number of frames. These two modalities are then processed with a dual-branch transformer model as shown in Figure 2, which outputs their counterparts $\hat{s}^{\alpha},\hat{s}^{\beta}\in\mathbb{R}^{N\times d_{s}}$, where $d_{s}$ denotes feature channels. After achieving features for each modality, we interact with them with cross attention, leading to a style-related embedding $\hat{s}$ aware of both audio and visual information.
@@
$\sigma_{s}$ $=\mathrm{softmax}(W_{s}\hat{s})\cdot(\hat{s}^{T}-\mu_{s})^{2},$ (3)
@@
where $W_{s}\in\mathbb{R}^{1\times d_{s}}$ is a trainable parameter, $\mu_{s},\sigma_{s}$ denotes the mean and variance of the learned style prior $s$.
Compared with the naive style encoder used in StyleTalk (Ma et al. 2023), our proposed model mainly enjoys the following merits: (1) As mentioned in Sec. 1, the audio information is vital for extracting intrinsic style, while StyleTalk cannot handle such a modality. Moreover, audio typically contains primarily information about the spoken content, thus making it non-trivial to extract information that is complimentary to the visual information contained in video frames. In comparison with StyleTalk, we design a specific structure to handle these complex data, considering both visual and audio information, leading to stronger style embedding. (2) Since the emotion of speakers would change as the video frames go on, it is sufficient to represent the intrinsic style with a deterministic feature, i.e. the same way as in StyleTalk. Our method, on the other hand, learns a better sequential embedding, which helps us model the style prior as a Gaussian distribution that is more representative.
Style-Driven Diffusion Process
After learning the intrinsic style prior from style reference videos, we can then control the talking head generation process with such a condition. Specifically, we build a talking head generation model based on previous methods such as V-Express (Wang et al. 2024). These methods apply various techniques to pretrained Stable Diffusion (SD) for better quality, including ReferenceNet for referencing an image, Audio Projection for embedding audio information, and temporal attention layer to enhance temporal coherence. In addition to these existing methods, we further propose two novel modules named HEAD-Kps Guider and Style Projection as follows that can better facilitate the input data.
@@
Each HEAD-Kps image spatially corresponds to the respective target frame. To fully utilize this, we use the HEAD-Kps Guider to encode the HEAD-Kps images. The Head-Kps images are constructed using landmarks from the upper half of the face, consisting of 8 keypoints. The HEAD-Kps Guider is a lightweight convolutional model that encodes each HEAD-Kps image into HEAD-Kps features, which match the shape of the latent. Subsequently, before being input into the denoising U-Net, the multi-frame latent features are directly added to the corresponding encoded HEAD-Kps features.
@@
To utilize the intrinsic style prior calculated in Sec. 3.1 to guide the denoising process, we first resample a style prior $s$ from the learned Gaussian distribution. Then $s$ is injected into the diffusion UNet through an additional style attention layer along with the specific cross attention layers for audio and temporal information.
Training Strategies
@@
Essentially, codes with similar intrinsic styles should cluster together in the style space. Therefore we apply contrastive learning to the style priors by constructing positive pairs $(s,s^{p})$ with the same identity and emotion, and negative pairs $(s,s^{n})$ with different identities or emotions. Then, the InfoNCE loss (Chen et al. 2020) with similarity metric $\zeta$ is enhanced between positive and negative sample pairs:
@@
where $\tau$ denotes a temperature parameter, $\mathcal{S}^{n}$ denotes all negative samples for $s$, and the similarity $\zeta(s_{i},s_{j})=\frac{1}{\|s_{i},s_{j}\|_{2}+1}\in(0,1]$ is an improved version obtained as the inverse of the $\mathcal{L}_{2}$ distance between sample pairs. We additionally add a fixed constant to stabilize the numerical range of the similarity and make the training process more stable.
@@
The training of the diffusion model adopts a three-stage progressive training method to gradually improve the model’s generative capability and stability, with the noise prediction loss as in Eq. 1 employed in each stage. (1) First, we train the model for single-frame image generation, where the diffusion UNet, ReferenceNet, and Head-Kps guider are involved in the training. In this stage, the diffusion UNet takes a single frame as input, the ReferenceNet processes different frames randomly selected from the same video clip, and the Head-Kps guider incorporates the encoded Head-Kps features into the latent space. Both the diffusion UNet and ReferenceNet initialize their weights from the original SD. (2) Second, we train the model for continuous multi-frame image generation, which includes the temporal module and the audio layer. In this stage, $f$ consecutive frames are sampled from a video clip and the parameters of ReferenceNet and Head-Kps guider are frozen. The temporal module initializes its weights from AnimateDiff (Guo et al. 2023). (3) After that the final stage is for transferring intrinsic style. In this stage, all other modules of the model are frozen, and only the style attention module is trained. This allows the model to generate corresponding facial expressions and details based on the intrinsic style input during the portrait image generation process.
Experiments
Figure 3: Visual comparison with recent SOTA methods. The first two rows show the comparison of reconstruction results, while the last two rows show the comparison of intrinsic style transfer results.

### v2

@@ Abstract
Project page: https://svportrait.github.io/
1 Introduction
Recent advancements in generative models have shed light on generating high-quality and realistic videos under various controlling conditions such as texts [17], images [2], videos [32], etc. Among all the different subtasks of video generation, Talking Head Generation (THG), as a human-centric task which aims to generate videos of talking heads guided by conditions such as speech and images, has emerged as a significant problem due to its wide application in scenarios such as digital humans, film production, virtual reality. In spite of its importance, this is one of the most challenging tasks in video generation, resulted from its low tolerance to artifacts in general and its demand of high fidelity in lip shapes, facial expressions, and head motions.
Following the commonly used generative models, the GAN-based THG methods [15, 36] have achieved remarkable results in generating high-resolution videos through adversarial training between generators and discriminators, particularly in terms of visual quality and lip-sync accuracy. Diffusion model-based THG methods [19, 21], on the other hand, excel in generating high-quality and high-resolution images and videos, and it outperforms GANs in terms of the stability and consistency of the generated content, thus becoming the mainstream methods for THG. These methods largely facilitate THG by strengthening the explicit controlling conditions such as facial keypoints and head motion sequences. However, they generally ignore the important fact of talking head videos. Essentially, when different people present speeches in real-life cases, they could have significant differences in habits and emotions under various circumstances. Such a fact in turn leads to different attributes in the corresponding talking head videos, including the visemes and expressions. Consequently, these habits and emotions are embedded as the intrinsic style in talking head videos. This intrinsic style, while being highly related to whether a video is realistic, can hardly be inferred from conditions such as facial keypoints which are widely adopted by the previous methods. As a result, when there is a large gap in the intrinsic style between the reference face and the speaker of the style reference video, the previous methods struggle to reproduce the real situation accurately.
@@ 1 Introduction
Specifically, our $\mathtt{SVP}$ focuses on two main problems, i.e. extracting intrinsic style embeddings from style reference videos and controlling diffusion models with such embeddings. For intrinsic style extraction, a naive solution would be following StyleTalk [13], which maps 3D Morphable Model (3DMM) [1] expression coefficients of the style reference video to style-related features. However, since attributes like visemes and expressions vary along the video frames, the deterministic embedding would suffer from insufficient capacity to model the latent manifold of intrinsic styles. Moreover, as one of the main parts of the video, the use of corresponding audio, which contains abundant information regarding the intrinsic styles, was not explored in StyleTalk, leading to unrepresentative style embeddings.
@@
To validate the effectiveness of our proposed method, we conduct extensive experiments and comparisons on the MEAD [27] and HDTF [37] datasets. Our method significantly outperforms other competitors across multiple metrics, including FVD [22], FID [8], PSNR, SSIM, the offset and confidence of SyncNet [6] and StyleSim. In addition to quantitative evaluation, we also perform comprehensive qualitative assessments. The results indicate that our method can generate highly natural and expressive talking videos, and can produce different emotions or even multiple changes in expressions within the same video according to user needs, achieving satisfactory visual effects.
@@
  • We propose an intrinsic style extractor that captures and expresses the intrinsic style of individuals via a self-supervised approach. We also incorporate audio information as an auxiliary into the style extractor to enhance the intrinsic style features. This allows the model to reflect the emotions and habits of speakers more accurately.
  • We designed a probabilistic style prior learning to adapt diffusion models. During the training of the style layer in the diffusion model, we sample the intrinsic style prior from the learned Gaussian distribution, enhancing stability and generalization capability.
2 Related Work
@@
There has been significant research on GAN-based methods for person-generic audio-driven talking head generation. Early methods [15, 5, 26] achieved lip synchronization by establishing a discriminator that correlates audio with lip movements. Other approaches [39, 38, 28, 29, 36] generated portrait videos by mapping audio to key facial information, such as landmarks, key points, or 3D Morphable Model (3DMM) [1] coefficients, before rendering the final frame. However, due to the limitations of GANs in terms of generative capacity, the results produced by these methods often suffer from artifacts like pseudo-textures or restricted motion ranges.
@@
Recently, there has been a surge of research  [19, 33, 21, 25, 30, 34, 35] utilizing diffusion models to achieve high-quality portrait videos. Among these, X-Portrait [33] and MegActor [35] rely on the pose and expression from the source video to generate the target video, which limits their ability to produce videos based solely on audio. DiffTalk [19] was the first to modify lip movements using audio and diffusion models, but it does not extend to driving other head parts. EMO [21] was the first to leverage LDM [17] and audio features to achieve overall motion in portraits. V-Express [25] controls the overall motion amplitude by adjusting audio attention weights, while Hallo [34] designed a hierarchical module to regulate the motion amplitude of different regions. In summary, current audio-driven diffusion model approaches have not taken into account that each portrait should exhibit a corresponding style while speaking, which is essential for generating higher-quality portrait videos.
@@
Previous research has explored several GAN-based methods  [27, 31, 9, 10, 12, 20, 13] for extracting style information to apply in talking head generation. MEAD [27] and Emotion [20] directly inject style labels into the network to drive the corresponding emotions. GC-AVT [12] and EAMM [10] map the facial expressions of each frame in the source video to each frame in the target video. LSF [31] and StyleTalk [13] employ 3D Morphable Models (3DMM) to extract facial information and construct style codes that drive the desired styles. Building on these approaches, our framework is the first to propose the extraction of intrinsic style priors by integrating audio and 3D facial information. Additionally, we are first to design a probabilistic learning to enhance style control within diffusion models.
3 Method
@@
The goal of THG is to generate a talking head video under the control of a reference portrait image, audio, Head-Kps image sequence, and intrinsic style prior. Among these conditions, the reference portrait image provides the background and facial identity, the audio guides the lip movements, and each Head-Kps image controls the head position and pose for each generated frame. The Head-Kps images are synthesized by mapping 8 facial keypoints onto a black background. These 8 facial keypoints correspond to the left and right edges of the face, the pupils of both eyes, and the bridge of the nose, which are used to guide the overall head movement. In addition, we propose probabilistic style prior learning to extract the style prior from the visual and audio content of a style reference video, which is used to determine facial emotion and speaking habit.
@@
In $\mathtt{SVP}$, we employ a Latent Diffusion Model (LDM) [17] to generate video frames. The LDM uses a diffusion and denoising process in the latent space via a Variational Autoencoder (VAE). It maps the input image $x$ to the latent space, encoding the image as $z=E(x)$, which helps maintain visual quality while reducing computational cost. During the diffusion process, Gaussian noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is gradually introduced into the latent $z$, degrading it into complete noise $z\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ after $T$ steps. In the reverse denoising process, the target latent $z$ is iteratively denoised from the sampled Gaussian noise using the diffusion model and then decoded by the VAE decoder $D$ into the output image $x=D(z)$. During training, given the latent $z_{0}=E(x_{0})$ and condition $c$, the denoising loss is:
@@
Among them, $z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$ represents the noisy latent variables at timestep $t\in[1,T]$, and $\epsilon_{t}$ is the added noise. $\epsilon_{\theta}$ is the noise predicted by the UNet model, modified using an attention mechanism with parameters $\theta$. This model employs a cross-attention mechanism to fuse the condition $c$ with the latent features $z_{t}$ , thereby guiding the image generation. $\mathtt{SVP}$ uses Stable Diffusion v1.5 (SDv1.5), a text-to-image Latent Diffusion Model (LDM), as the backbone. SDv1.5 is implemented based on U-Net [18], with each Transformer [24] block containing both self-attention and cross-attention layers.
@@
As depicted in Figure 2, Our $\mathtt{SVP}$ consists of two important designs, namely Probabilistic Style Prior Learning and Style-Driven Diffusion Process. In Probabilistic Style Prior Learning, we propose the style extractor takes the 3DMM expression parameters and audio features as inputs to generate the style prior, which is represented as a Gaussian distribution. In the style-driven diffusion process, the encoded Head-Kps sequence, reference image, audio features, and style prior are progressively input into the Denoising UNet as control conditions through their respective attention layers. Finally, the vivid portrait frames are generated by the VAE decoder after the iteration of the denoising process.
3.1 Probabilistic Style Prior Learning
In order to learn representative intrinsic style indicators from style reference videos, we propose the novel probabilistic style prior learning. Built upon the transformer-based style encoder as in StyleTalk [13], we adopt a novel framework to make better usage of the style-related information contained in each video. Concretely, for a video clip, we first transform it into its corresponding frame-level audio parameters $\alpha\in\mathbb{R}^{N\times 1920}$ via Whisper-Tiny [16] and sequential expression parameters $\beta\in{\mathbb{R}}^{N\times 64}$ via the 3DMM encoder, where $N$ denotes number of frames. These two modalities are then processed with a dual-branch transformer model as shown in Figure 2, which outputs their counterparts $\hat{s}^{\alpha},\hat{s}^{\beta}\in\mathbb{R}^{N\times d_{s}}$, where $d_{s}$ denotes feature channels. After achieving features for each modality, we interact with them with cross attention, leading to a style-related embedding $\hat{s}$ aware of both audio and visual information.
@@ 3.1 Probabilistic Style Prior Learning
$\sigma_{s}^{2}$ $=\mathrm{softmax}(W_{s}\hat{s})\cdot(\hat{s}^{T}-\mu_{s})^{2},$ (3)
@@
where $W_{s}\in\mathbb{R}^{1\times d_{s}}$ is a trainable parameter, $\mu_{s},\sigma_{s}^{2}$ denotes the mean and variance of the learned style prior $s$.
Compared with the naive style encoder used in StyleTalk [13], our proposed model mainly enjoys the following merits: (1) As mentioned in Sec. 1, the audio information is vital for extracting intrinsic style, while StyleTalk cannot handle such a modality. Moreover, audio typically contains primarily information about the spoken content, thus making it non-trivial to extract information that is complimentary to the visual information contained in video frames. In comparison with StyleTalk, we design a specific structure to handle these complex data, considering both visual and audio information, leading to stronger style embedding. (2) Since the emotion of speakers would change as the video frames go on, it is sufficient to represent the intrinsic style with a deterministic feature, i.e. the same way as in StyleTalk. Our method, on the other hand, learns a better sequential embedding, which helps us model the style prior as a Gaussian distribution that is more representative.
3.2 Style-Driven Diffusion Process
After learning intrinsic style from style reference videos, we can control details such as facial expressions with such a condition, enabling a more refined talking head generation process. Specifically, we build a talking head generation model based on previous methods such as V-Express [25]. These methods apply various techniques to pretrained Stable Diffusion (SD) for better quality. For instance, ReferenceNet can generate similar feature maps and integrate the extracted features into the diffusion backbone, preserving the visual information of the face and background from the reference image. The audio projection module, which embeds audio information, controls the generation of lip movements through a cross-attention mechanism. The temporal attention layer, which enhances temporal coherence by performing self-attention on the frame sequence to capture inter-frame correlations. In addition to these existing methods, we further propose two novel modules named HEAD-Kps Guider and Style Projection as follows that can better facilitate the input data.
@@ 3.2 Style-Driven Diffusion Process
Each HEAD-Kps image spatially corresponds to the respective target frame, containing information about the head’s position and rotation. To fully utilize this, we use the HEAD-Kps Guider to encode the HEAD-Kps images. The Head-Kps images are constructed using landmarks from the upper half of the face, consisting of 8 keypoints. The HEAD-Kps Guider is a lightweight convolutional model that encodes the keypoints into HEAD-Kps features, which represent spatial information and match the shape of the latent features. Subsequently, before being input into the denoising U-Net, the multi-frame latent features are directly added to the corresponding encoded HEAD-Kps features, enabling the model to accurately interpret the head’s spatial information.
@@
To utilize the intrinsic style priors obtained in Sec. 3.1 to guide the denoising process, we first resample a corresponding intrinsic style prior $s$ from the Gaussian distribution learned from the style reference video. Then $s$ is injected into the diffusion UNet through an additional style attention layer, where it interacts with other features via a cross-attention mechanism to supplement additional facial details such as expressions and speaking habits. The intrinsic style prior information can be injected into the spatial cross-attention layer to provide spatial knowledge as follows:
$z_{s}=z_{a}+\mathrm{CrossAttn}\left(Q(z_{a}),K(s),V(s)\right),$ (5)
where $z_{a}$ is the spatial latent features after being injected with reference attention and audio attention, and $z_{s}$ is the adjusted spatial features guided by intrinsic style prior spatial-aware level.
3.3 Training Strategies
@@
Essentially, codes with similar intrinsic styles should cluster together in the style space. Therefore we apply contrastive learning to the style priors by constructing positive pairs $(s,s^{p})$ with the same identity and emotion, and negative pairs $(s,s^{n})$ with different identities or emotions. Then, the InfoNCE loss [4] with similarity metric $\zeta$ is enhanced between positive and negative sample pairs:
@@
where $\tau$ denotes a temperature parameter, $\mathcal{S}^{n}$ denotes all negative samples for $s$, and the similarity $\zeta(s_{i},s_{j})=\frac{1}{\|s_{i}-s_{j}\|_{2}+1}\in(0,1]$ is an improved version obtained as the inverse of the $\mathcal{L}_{2}$ distance between sample pairs. We additionally add a fixed constant to stabilize the numerical range of the similarity and make the training process more stable.
@@
The training of the diffusion model adopts a three-stage progressive training method to gradually improve the model’s generative capability and stability, with the noise prediction loss as in Eq. 1 employed in each stage. (1) First, we train the model for single-frame image generation, where the diffusion UNet, ReferenceNet, and Head-Kps guider are involved in the training. In this stage, the diffusion UNet takes a single frame as input, the ReferenceNet processes different frames randomly selected from the same video clip, and the Head-Kps guider incorporates the encoded Head-Kps features into the latent space. Both the diffusion UNet and ReferenceNet initialize their weights from the original SD. (2) Second, we train the model for continuous multi-frame image generation, which includes the temporal module and the audio layer. In this stage, $f$ consecutive frames are sampled from a video clip and the parameters of ReferenceNet and Head-Kps guider are frozen. The temporal module initializes its weights from AnimateDiff [7]. (3) After that the final stage is for transferring intrinsic style. In this stage, all other modules of the model are frozen, and only the style attention module is trained. This allows the model to generate corresponding facial expressions and details based on the intrinsic style input during the portrait image generation process.
4 Experiments
Figure 3: (a) Visual comparison with recent SOTA methods. The first two rows show the comparison of reconstruction results, while the last two rows show the comparison of intrinsic style transfer results. (b) Our method uses intrinsic style to generate frames on different types of portraits. This demonstrates that our method can successfully apply intrinsic style to various types of data, even if only real-life videos are available for training.